# Interconnect-Aware Logic Resynthesis for Multi-Die FPGAs

### v1

@@ 1. Introduction
Interposer-based multi-die integration has emerged as a practical solution to overcome the reticle size limits of lithography, enabling large systems to be implemented by partitioning designs across multiple dies (Radojcic, 2017). This approach is now widely adopted in both application-specific integrated circuits (ASICs) and field-programmable gate arrays (FPGAs(Graening et al., 2025; Hahn Pereira and Betz, 2014). However, distributing a design across dies inevitably introduces inter-die connections, i.e., links that cross boundaries between different dies, which are also referred to as super-long lines (SLLs).
Compared to intra-die connections that remain within a single die, inter-die connections raise two major challenges. First, they incur significantly larger delay because the signals need to traverse long physical distances across dies through the interposer and pass through additional interface circuitry. Thus, timing paths including inter-die connections are much slower, and even a small number of such connections can noticeably increase the critical path delay This effect is particularly profound in multi-die FPGAs, where signals can cross die boundaries only through dedicated inter-die links, making timing closure more difficult to achieve (Boutros et al., 2023; Nasiri et al., 2016; Hahn Pereira and Betz, 2014). Second, excessive inter-die connections increase physical design complexity. Inter-die connections consume scarce die-to-die links, which constrain block placement and tighten routing constraints. Prior work on single-die FPGAs has shown that higher interconnect complexity leads to longer physical design runtime (Wang and Stroobandt, 2025). In multi-die CAD flows, this effect can be further amplified. Inter-die connections tightly couple implementations on different dies, so physical implementation decisions on one die affect others. This coupling limits independent processing of dies, reduces parallelism in the design flow, and increases overall implementation difficulty and runtime.
To address these issues, prior work has explored reducing inter-die connections for FPGAs at different stages of the CAD flow. These efforts include multi-die-aware placement (Di et al., 2024) and co-optimization of high-level synthesis directives with floorplanning (Guo et al., 2021; Du et al., 2023), highlighting that inter-die delay is a critical concern across different abstraction levels. Among them, partitioning-based FPGA CAD flows reduce inter-die connections through timing-aware assignment of blocks across dies (Kuo et al., 2018; Liao and Mak, 2020). Previous work integrates timing-driven partitioning into the open-source Verilog-to-Routing (VTR) framework (Luu et al., 2014), demonstrating improvements in critical-path delay and maximum operating frequency by discouraging cuts on timing-critical paths (Iyer et al., 2025; Raikar and Stroobandt, 2024).
@@ 2. Preliminaries
Figure 1. Lateral and top views of a four-die FPGA system (AMD Xilinx Alveo U250 (AMD Xilinx, 2023)).
@@ 2.1. Multi-Die FPGAs
Commercial multi-die FPGAs may further employ tile-based interposer interfaces with dedicated interface tiles (e.g., Laguna columns) (Ravishankar et al., 2018) shown in Figure 2. Since the logic resynthesis techniques studied in this work operate on a generic logic netlist representation, they are agnostic to the specific implementation of the interposer. For clarity and compatibility with open-source academic CAD tools, we adopt the routing-based abstraction throughout this paper.
Table 1 summarizes representative delay values for intra-die routing wires and inter-die interposer links used in this work. The percentage indicates the fraction of routing tracks of each wire type within the total routing channel width. These values are derived from published technology-scaled delay models (Petelin and Betz, 2016; Sun et al., 2019) and serve as parameters for characterizing the timing differences between routing resources.
@@ 2.2. Partitioning-Based Multi-Die FPGA Flow
By decomposing a monolithic design into multiple sub-circuits on multiple dies, partitioning reduces the optimization space of the CAD problems and enables scalable and partially parallel CAD processing, while still requiring careful coordination of inter-die connections. Figure 3 illustrates two representative partitioning-based CAD strategies for multi-die FPGA implementation, which differ in whether partitioning is performed before or after packing. In packing-first flows, LUTs are first clustered into architecture-specific blocks, such as configurable logic blocks (CLBs), and partitioning is subsequently applied to these packed blocks (Iyer et al., 2025; Graening et al., 2025). This ordering closely reflects physical constraints and simplifies downstream placement and routing, but fixes the logic connectivity at the granularity of packed blocks. In contrast, partitioning-first flows partition the primitive netlist before packing, allowing logic connectivity to be considered before architectural constraints are fully imposed. Several state-of-art multi-die CAD flows (Iyer et al., 2025; Graening et al., 2025; Raikar and Stroobandt, 2023, 2024) use the partitioning-first paradigm. The netlist is first partitioned into sub-circuits on multiple dies, after which packing, placement, and routing are performed for each die.
@@ 3.1. Motivation and Overview
This work focuses on reducing the number of SLLs by targeting the logic-level origins of inter-die connectivity. These inter-die connections are not purely required by architectural constraints, but often stem from the logic implementation of the mapped LUT-level netlist. Consequently, modifying the logic implementation can often replace cross-die signals with functionally equivalent in-die implementations, thereby reducing the inter-die connections. By directly restructuring the logic to reduce inter-die connections, ResynMD tool flow complements existing partitioning-first CAD flows for multi-die FPGAs (Raikar and Stroobandt, 2024; Di et al., 2024) (see Figure 3(b)), which optimize inter-die connections through partitioning or placement without altering the circuit at the logic level.
@@
Compared to the existing partitioning-first CAD flows, such as LiquidMD (Raikar and Stroobandt, 2024; Vercruyce et al., 2017, 2018), the key distinction of the proposed tool flow in Figure 4 is the SLL-aware logic resynthesis step that is applied after partitioning but before physical implementation, as illustrated by the blue segment in Figure 4. To the best of our knowledge, ResynMD is the first approach that explicitly targets logic-level restructuring for reducing inter-die connections. This introduces a new optimization dimension that is complementary to existing partitioning-based and placement-based multi-die optimization techniques and can provide synergistic benefits when integrated into current multi-die FPGA design flows. The following subsection details the problem formulation and the core components of the proposed SLL-aware resynthesis method.
@@ 3.2. Problem Formulation of SLL-Aware Logic Resynthesis
Conventional logic synthesis methods, such as  (Mishchenko and Brayton, 2006; Lee and others, 2021; Amarú et al., 2018; Liu and others, 2024), optimize metrics such as area or delay without considering the multi-die context and do not explicitly reduce inter-die connections. In contrast, the proposed SLL-aware logic resynthesis operates on a partitioned LUT-level netlist and directly targets inter-die connections. In this setting, the die assignment of each LUT is fixed by the partitioning stage, and only functionally equivalent logic transformations are permitted.
@@
Because global optimizing Equation (2) over all equivalent netlists needs exploring a huge combinatorial search space, we adopt a greedy heuristic that iteratively applies local resubstitution (Mishchenko and Brayton, 2006) transformations, which preserve functionality while modifying the fanin structure of nodes. This approach offers good scalability and computational efficiency for large-scale circuits. In each iteration, a transformation is accepted only if it strictly reduces the total number of SLLs and satisfies the area constraint. More details on the resubstitution procedure are presented in Section 3.3.
@@ 12     New node $v^{\prime}\leftarrow LUTImplement(\hat{f}_{v})$;
This subsection presents the core algorithm of the proposed SLL-aware logic resynthesis method. The method builds upon the resynthesis technique MFS (Mishchenko et al., 2011). Unlike conventional MFS, which optimizes traditional metrics such as area or delay, our new contributions lie in adapting the resynthesis approach to the multi-die context with a specific focus on reducing inter-die connections, as illustrated in Figure 5. The inter-die connections are reduced by performing LUT-level resubstitution. Here, resubstitution (Mishchenko et al., 2011, 2006) refers to a local logic transformation that replaces a node with a functionally equivalent implementation expressed in terms of a selected set of other nodes (called divisors) in the netlist.
@@ 3.3.2. Window Construction
This step corresponds to Line 1 in Algorithm 1. We adopt the same window construction strategy as MFS (Mishchenko et al., 2011). As illustrated in the middle of Figure 5(b), for each pivot node $v$ (i.e., node being processed for resubstitution), its window $W(v)$ is a sub-circuit surrounding $v$ that captures the local logic context relevant for $v$’s resubstitution. $W(v)$ is formed by expanding its TFOs up to a limited depth $D_{1}$. Then, $D_{2}$ levels of TFIs of the pivot node $v$ determines the window primary inputs (PIs). After that, the internal nodes are obtained through a topologically ordered traversal of the induced subgraph, forming the window $W(v)$.
@@ 13  $S_{d}^{\prime}\leftarrow BaseDivisors\cup\{d\}$;
For each constructed divisors subset $S_{d}^{\prime}$, a SAT-based existence check (Mishchenko et al., 2011) is performed under the extracted care set $C$ to determine whether $f_{v}$ can be equivalently expressed using signals in $S_{d}^{\prime}$. If the SAT-based existence check succeeds, an interpolation is extracted to construct the equivalent function $\hat{f}_{v}$ (Lee et al., 2007). The interpolation depends only on signals in $S_{d}^{\prime}$ and preserves equivalence under the care set $C$. The derived function $\hat{f}_{v}$ is then implemented as a new LUT node $v^{\prime}$ whose fanins correspond to the selected divisors in $S_{d}^{\prime}$. Node $v^{\prime}$ is inserted into the original network, and all fanout connections of the original pivot node $v$ are redirected to $v^{\prime}$. Subsequently, node $v$ and any newly unreachable logic in its $\mathrm{MFFC}(v)$ are removed from the network. This network update corresponds to line 1 in Algorithm 1.
@@ 3.4. Partition Imbalance Induced by Inter-Die Resynthesis
In addition to minimizing the number of inter-die connections, partitioning-based multi-die FPGA CAD flows are also expected to maintain a balanced distribution of logic resources across dies. Partition balance is a key feasibility constraint: a severely imbalanced die can become overutilized, leading to packing/placement failures or degraded QoR. A commonly used metric is the partition imbalance ratio, denoted by $\rho$, which quantifies the deviation of logic distribution from an ideal uniform partition (Raikar and Stroobandt, 2023). It is defined as
@@ 4.1. Experimental Setup
Our interconnect-aware logic resynthesis method for multi-die FPGAs, ResynMD, is implemented in C++ and Java. All experiments are conducted on a workstation equipped with an AMD Ryzen 9 7900X 12-core processor and 128 GB RAM, running a 64-bit Linux system. For the logic synthesis and technology mapping stage in Figure 4, we use the Yosys (Wolf, 2025) and ABC (Brayton and Mishchenko, 2010) tools. For the partitioning stage, we utilize an hMETIS-based (Karypis and Kumar, 1998) timing-aware partitioning tool. For the physical implementation stage, we utilize the AAPack (Luu et al., 2011) clustering algorithm provided in VTR8 (Murray et al., 2020), and the multi-die placement algorithm implemented in LiquidMD (Raikar and Stroobandt, 2024; Vercruyce et al., 2017). No post-routing results are reported in this work, while our tool flow can be extended to evaluate routing performance in future work.
We fix the imbalance factor $UB$ of the partitioner to 1.25 across all experiments. We evaluate our approach using three benchmark suites, each targeting a distinct evaluation objective. The EPFL (Amarú et al., 2015) benchmark suite consists of purely combinational designs and is used to quantify the intrinsic effectiveness of the proposed logic-level resynthesis in reducing inter-die connections. Since these circuits are relatively small and do not reflect realistic physical-design constraints, we report only connectivity-level metrics on EPFL benchmarks.
To evaluate physical-design impact in the presence of sequential logic, we additionally use a set of MCNC (Yang, 1991) benchmarks. These circuits contain flip-flops but no heterogeneous hard blocks, which enables isolating the effect of sequential elements while still allowing post-placement evaluation.
To evaluate physical-design impact on realistic heterogeneous FPGA designs, we utilize the Koios (Arora et al., 2021) benchmark suite, which comprises designs featuring hard blocks such as DSPs and block RAMs. In these circuits, hard macros and timing-optimized sequential elements are preserved and treated as structural boundaries, allowing us to assess whether SLL-aware resynthesis remains beneficial under constrained rewriting flexibility. For both MCNC and Koios, we report post-placement results to quantify downstream physical-design effects.
@@ 4.2. Multi-Die Cost Modeling and Connectivity Metrics
During placement, wirelength is commonly estimated by summing the contributions of all nets, where the wirelength of each net is approximated using the half-perimeter wirelength (HPWL) of its bounding box. To account for net-specific characteristics, the HPWL of each net is multiplied by a weighting factor $q(n)$ (Vercruyce et al., 2017; Raikar and Stroobandt, 2024). The overall wirelength cost is thus defined as
@@
Figure 7. Bounding-box ($B$) wirelength estimation before and after die partitioning (Raikar and Stroobandt, 2023). (a) A non-partitioned net enclosed by a single bounding box within one die. (b) After partitioning, the net spans multiple dies, forming die-local terminal clusters. (c) The partitioned net is modeled using multiple die-local bounding boxes connected by inter-die segments, and the total $B$ wirelength accounts for both intra-die and inter-die components.

### v2

@@ 1. Introduction
Interposer-based multi-die integration has emerged as a practical solution to overcome the reticle size limits of lithography, enabling large systems to be implemented by partitioning designs across multiple dies (radojcic2017more, ). This approach is now widely adopted in both application-specific integrated circuits (ASICs) and field-programmable gate arrays (FPGAs(graening2025chipletpartcostawarepartitioning25d, ; Hahn2014fpga, ). However, distributing a design across dies inevitably introduces inter-die connections, i.e., links that cross boundaries between different dies, which are also referred to as super-long lines (SLLs).
Compared to intra-die connections that remain within a single die, inter-die connections raise two major challenges. First, they incur significantly larger delay because the signals need to traverse long physical distances across dies through the interposer and pass through additional interface circuitry. Thus, timing paths including inter-die connections are much slower, and even a small number of such connections can noticeably increase the critical path delay This effect is particularly profound in multi-die FPGAs, where signals can cross die boundaries only through dedicated inter-die links, making timing closure more difficult to achieve (boutros2023into, ; Nasiri2016TVLSI, ; Hahn2014fpga, ). Second, excessive inter-die connections increase physical design complexity. Inter-die connections consume scarce die-to-die links, which constrain block placement and tighten routing constraints. Prior work on single-die FPGAs has shown that higher interconnect complexity leads to longer physical design runtime (wang2024slip, ). In multi-die CAD flows, this effect can be further amplified. Inter-die connections tightly couple implementations on different dies, so physical implementation decisions on one die affect others. This coupling limits independent processing of dies, reduces parallelism in the design flow, and increases overall implementation difficulty and runtime.
To address these issues, prior work has explored reducing inter-die connections for FPGAs at different stages of the CAD flow. These efforts include multi-die-aware placement (LEAPS2024TCASI, ) and co-optimization of high-level synthesis directives with floorplanning (guo2021autobridge_fccm, ; FACO2023FPGA, ), highlighting that inter-die delay is a critical concern across different abstraction levels. Among them, partitioning-based FPGA CAD flows reduce inter-die connections through timing-aware assignment of blocks across dies (kuo2018pin, ; liao2020pin, ). Previous work integrates timing-driven partitioning into the open-source Verilog-to-Routing (VTR) framework (jasonluu2017vtr7, ), demonstrating improvements in critical-path delay and maximum operating frequency by discouraging cuts on timing-critical paths (iyer_partitioning-based_2025, ; raveena2024HEART, ).
@@ 2. Preliminaries
Figure 1. Lateral and top views of a four-die FPGA system (AMD Xilinx Alveo U250 (amd-u200-ds962, )).
@@ 2.1. Multi-Die FPGAs
Commercial multi-die FPGAs may further employ tile-based interposer interfaces with dedicated interface tiles (e.g., Laguna columns) (ravishankar2018sat, ) shown in Figure 2. Since the logic resynthesis techniques studied in this work operate on a generic logic netlist representation, they are agnostic to the specific implementation of the interposer. For clarity and compatibility with open-source academic CAD tools, we adopt the routing-based abstraction throughout this paper.
Table 1 summarizes representative delay values for intra-die routing wires and inter-die interposer links used in this work. The percentage indicates the fraction of routing tracks of each wire type within the total routing channel width. These values are derived from published technology-scaled delay models (Oleg2016Delay, ; Sun2019Pattern, ) and serve as parameters for characterizing the timing differences between routing resources.
@@ 2.2. Partitioning-Based Multi-Die FPGA Flow
By decomposing a monolithic design into multiple sub-circuits on multiple dies, partitioning reduces the optimization space of the CAD problems and enables scalable and partially parallel CAD processing, while still requiring careful coordination of inter-die connections. Figure 3 illustrates two representative partitioning-based CAD strategies for multi-die FPGA implementation, which differ in whether partitioning is performed before or after packing. In packing-first flows, LUTs are first clustered into architecture-specific blocks, such as configurable logic blocks (CLBs), and partitioning is subsequently applied to these packed blocks (iyer_partitioning-based_2025, ; graening2025chipletpartcostawarepartitioning25d, ). This ordering closely reflects physical constraints and simplifies downstream placement and routing, but fixes the logic connectivity at the granularity of packed blocks. In contrast, partitioning-first flows partition the primitive netlist before packing, allowing logic connectivity to be considered before architectural constraints are fully imposed. Several state-of-art multi-die CAD flows (iyer_partitioning-based_2025, ; graening2025chipletpartcostawarepartitioning25d, ; raveena2022slip, ; raveena2024HEART, ) use the partitioning-first paradigm. The netlist is first partitioned into sub-circuits on multiple dies, after which packing, placement, and routing are performed for each die.
@@ 3.1. Motivation and Overview
This work focuses on reducing the number of SLLs by targeting the logic-level origins of inter-die connectivity. These inter-die connections are not purely required by architectural constraints, but often stem from the logic implementation of the mapped LUT-level netlist. Consequently, modifying the logic implementation can often replace cross-die signals with functionally equivalent in-die implementations, thereby reducing the inter-die connections. By directly restructuring the logic to reduce inter-die connections, ResynMD tool flow complements existing partitioning-first CAD flows for multi-die FPGAs (raveena2024HEART, ; LEAPS2024TCASI, ) (see Figure 3(b)), which optimize inter-die connections through partitioning or placement without altering the circuit at the logic level.
@@
Compared to the existing partitioning-first CAD flows, such as LiquidMD (raveena2024HEART, ; Dries2017LiquidFPT, ; Dries2018MultiPart, ), the key distinction of the proposed tool flow in Figure 4 is the SLL-aware logic resynthesis step that is applied after partitioning but before physical implementation, as illustrated by the blue segment in Figure 4. To the best of our knowledge, ResynMD is the first approach that explicitly targets logic-level restructuring for reducing inter-die connections. This introduces a new optimization dimension that is complementary to existing partitioning-based and placement-based multi-die optimization techniques and can provide synergistic benefits when integrated into current multi-die FPGA design flows. The following subsection details the problem formulation and the core components of the proposed SLL-aware resynthesis method.
@@ 3.2. Problem Formulation of SLL-Aware Logic Resynthesis
Conventional logic synthesis methods, such as  (mishchenko2006scalable, ; lee2021simulation, ; amaru2018improvements, ; liu2024cbtune, ), optimize metrics such as area or delay without considering the multi-die context and do not explicitly reduce inter-die connections. In contrast, the proposed SLL-aware logic resynthesis operates on a partitioned LUT-level netlist and directly targets inter-die connections. In this setting, the die assignment of each LUT is fixed by the partitioning stage, and only functionally equivalent logic transformations are permitted.
@@
Because global optimizing Equation (2) over all equivalent netlists needs exploring a huge combinatorial search space, we adopt a greedy heuristic that iteratively applies local resubstitution (mishchenko2006scalable, ) transformations, which preserve functionality while modifying the fanin structure of nodes. This approach offers good scalability and computational efficiency for large-scale circuits. In each iteration, a transformation is accepted only if it strictly reduces the total number of SLLs and satisfies the area constraint. More details on the resubstitution procedure are presented in Section 3.3.
@@ 12     New node $v^{\prime}\leftarrow LUTImplement(\hat{f}_{v})$;
This subsection presents the core algorithm of the proposed SLL-aware logic resynthesis method. The method builds upon the resynthesis technique MFS (alanmi2011mfs, ). Unlike conventional MFS, which optimizes traditional metrics such as area or delay, our new contributions lie in adapting the resynthesis approach to the multi-die context with a specific focus on reducing inter-die connections, as illustrated in Figure 5. The inter-die connections are reduced by performing LUT-level resubstitution. Here, resubstitution (alanmi2011mfs, ; mishchenko2006using, ) refers to a local logic transformation that replaces a node with a functionally equivalent implementation expressed in terms of a selected set of other nodes (called divisors) in the netlist.
@@ 3.3.2. Window Construction
This step corresponds to Line 1 in Algorithm 1. We adopt the same window construction strategy as MFS (alanmi2011mfs, ). As illustrated in the middle of Figure 5(b), for each pivot node $v$ (i.e., node being processed for resubstitution), its window $W(v)$ is a sub-circuit surrounding $v$ that captures the local logic context relevant for $v$’s resubstitution. $W(v)$ is formed by expanding its TFOs up to a limited depth $D_{1}$. Then, $D_{2}$ levels of TFIs of the pivot node $v$ determines the window primary inputs (PIs). After that, the internal nodes are obtained through a topologically ordered traversal of the induced subgraph, forming the window $W(v)$.
@@ 13  $S_{d}^{\prime}\leftarrow BaseDivisors\cup\{d\}$;
For each constructed divisors subset $S_{d}^{\prime}$, a SAT-based existence check (alanmi2011mfs, ) is performed under the extracted care set $C$ to determine whether $f_{v}$ can be equivalently expressed using signals in $S_{d}^{\prime}$. If the SAT-based existence check succeeds, an interpolation is extracted to construct the equivalent function $\hat{f}_{v}$ (lee2007interpolation, ). The interpolation depends only on signals in $S_{d}^{\prime}$ and preserves equivalence under the care set $C$. The derived function $\hat{f}_{v}$ is then implemented as a new LUT node $v^{\prime}$ whose fanins correspond to the selected divisors in $S_{d}^{\prime}$. Node $v^{\prime}$ is inserted into the original network, and all fanout connections of the original pivot node $v$ are redirected to $v^{\prime}$. Subsequently, node $v$ and any newly unreachable logic in its $\mathrm{MFFC}(v)$ are removed from the network. This network update corresponds to line 1 in Algorithm 1.
@@ 3.4. Partition Imbalance Induced by Inter-Die Resynthesis
In addition to minimizing the number of inter-die connections, partitioning-based multi-die FPGA CAD flows are also expected to maintain a balanced distribution of logic resources across dies. Partition balance is a key feasibility constraint: a severely imbalanced die can become overutilized, leading to packing/placement failures or degraded QoR. A commonly used metric is the partition imbalance ratio, denoted by $\rho$, which quantifies the deviation of logic distribution from an ideal uniform partition (raveena2022slip, ). It is defined as
@@ 4.1. Experimental Setup
Our interconnect-aware logic resynthesis method for multi-die FPGAs, ResynMD, is implemented in C++ and Java. All experiments are conducted on a workstation equipped with an AMD Ryzen 9 7900X 12-core processor and 128 GB RAM, running a 64-bit Linux system. For the logic synthesis and technology mapping stage in Figure 4, we use the Yosys (yosys2025yosys, ) and ABC (brayton2010abc, ) tools. For the partitioning stage, we utilize an hMETIS-based (karypis1998hypergraph, ) timing-aware partitioning tool. For the physical implementation stage, we utilize the AAPack (Luu2011AAPack, ) clustering algorithm provided in VTR8 (murray_vtr_2020, ), and the multi-die placement algorithm implemented in LiquidMD (raveena2024HEART, ; Dries2017LiquidFPT, ). No post-routing results are reported in this work, while our tool flow can be extended to evaluate routing performance in future work.
We fix the imbalance factor $UB$ of the partitioner to 1.25 across all experiments. We evaluate our approach using three benchmark suites, each targeting a distinct evaluation objective. The EPFL (epfl, ) benchmark suite consists of purely combinational designs and is used to quantify the intrinsic effectiveness of the proposed logic-level resynthesis in reducing inter-die connections. Since these circuits are relatively small and do not reflect realistic physical-design constraints, we report only connectivity-level metrics on EPFL benchmarks.
To evaluate physical-design impact in the presence of sequential logic, we additionally use a set of MCNC (yang1991logic, ) benchmarks. These circuits contain flip-flops but no heterogeneous hard blocks, which enables isolating the effect of sequential elements while still allowing post-placement evaluation.
To evaluate physical-design impact on realistic heterogeneous FPGA designs, we utilize the Koios (arora2021koios, ) benchmark suite, which comprises designs featuring hard blocks such as DSPs and block RAMs. In these circuits, hard macros and timing-optimized sequential elements are preserved and treated as structural boundaries, allowing us to assess whether SLL-aware resynthesis remains beneficial under constrained rewriting flexibility. For both MCNC and Koios, we report post-placement results to quantify downstream physical-design effects.
@@ 4.2. Multi-Die Cost Modeling and Connectivity Metrics
During placement, wirelength is commonly estimated by summing the contributions of all nets, where the wirelength of each net is approximated using the half-perimeter wirelength (HPWL) of its bounding box. To account for net-specific characteristics, the HPWL of each net is multiplied by a weighting factor $q(n)$ (Dries2017LiquidFPT, ; raveena2024HEART, ). The overall wirelength cost is thus defined as
@@
Figure 7. Bounding-box ($B$) wirelength estimation before and after die partitioning (raveena2022slip, ). (a) A non-partitioned net enclosed by a single bounding box within one die. (b) After partitioning, the net spans multiple dies, forming die-local terminal clusters. (c) The partitioned net is modeled using multiple die-local bounding boxes connected by inter-die segments, and the total $B$ wirelength accounts for both intra-die and inter-die components.